OpenStudy (anonymous):

Find the derivative of the function y defined implicitly in terms of x y^2 = ln(3x+1y)

OpenStudy (anonymous):

i'm confused about the explicitly in terms of x

OpenStudy (anonymous):

explicitly is in y= some equation with "x"

OpenStudy (anonymous):

implicitly you take the derivative of the entire equation

OpenStudy (anonymous):

\[\frac{d}{dx}[y^2] = 2y(\frac{dy}{dx})\] are you familiar with this?

OpenStudy (anonymous):

yes but i must have done it wrong I got 3/ (6xy-y+2y^2)

OpenStudy (anonymous):

\[\frac{d}{dx}[\ln(u)] = \frac{u'}{u}\] did you use this rule as well?

OpenStudy (anonymous):

for ln, i used 1/x times derivative of x

OpenStudy (anonymous):

oh i got it :) but thanks for clarifying the implicit in terms of y thing. i interpreted it as d/dx

OpenStudy (anonymous):

so how woudl you know it was d/dx or d/dy and what does that mean?

OpenStudy (anonymous):

yes, same thing, \[\frac{ 1 }{ x } * x' = \frac {x'}{x}\] let's see if i can do this. we got the left part: \[2y\frac{dy}{dx} = \frac{3 + \frac{dy}{dx}}{3x+y}\]

OpenStudy (anonymous):

d/dx is the derivative with respect to x and d/dy is the derivative with respect to y

OpenStudy (anonymous):

here's a few examples:

OpenStudy (anonymous):

\[\frac{d}{dx}[x] = 1\] \[\frac{d}{dx}[y] = \frac{dy}{dx}\] \[\frac{d}{dt}[x] = \frac{dx}{dt}\] \[\frac{d}{dt}[y] = \frac{dy}{dt}\]

OpenStudy (anonymous):

basically the d/d means "the derivative of" and "y" in the numerator means "the derivative of y" and an "x" in the denominator (dy/dx) means "the derivative of y with respect to x"

OpenStudy (anonymous):

what about dy/dx [y] or does that happen?

OpenStudy (anonymous):

dy/dx[y] = 0 because there is no X variables to respect. we assume all other variables are constant

OpenStudy (anonymous):

oh ok. Thank you!!

OpenStudy (anonymous):

i dont think "y" can be in the derivative operator but if it was y = h^2 and dy/dx[y] then it's 0

OpenStudy (anonymous):

since y is a function

OpenStudy (anonymous):

or equation or w/e

OpenStudy (raden):

y^2 = ln(3x+1y) 2y dy = 3 dx/(3x+1) dy/dx=3/(2y*(3x+1)) y'=3/(2y*(3x+1)) y'=3/(2sqrt(ln(3x+1))*(3x+1))

OpenStudy (anonymous):

thanks to you too!

OpenStudy (raden):

(3x+y) i mean * lol